# BEVPR: BEV-Enhanced Visual Place Recognition with Structural Cues

### v1

@@ Abstract
In this paper, we propose a new image-based visual place recognition (VPR) framework by exploiting the structural cues in bird’s-eye view (BEV) from a single monocular camera. The motivation arises from two key observations about VPR: 1) For the methods based on both camera and LiDAR sensors, the integration of LiDAR in robotic systems has led to increased expenses, while the alignment of data between different sensors is also a major challenge. 2) Other image-/camera-based methods, involving integrating RGB images and their derived variants (e.g., pseudo depth images, pseudo 3D point clouds), exhibit several limitations, such as the failure to effectively exploit the explicit spatial relationships between different objects. To tackle the above issues, we design a new BEV-enhanced VPR framework, namely BEVPR, which can generate a composite descriptor with both visual cues and spatial awareness solely based on a single camera. For the visual cues, any popular aggregation module for RGB global features can be integrated into our framework. The key points lie in: 1) We use BEV segmentation features as an explicit source of structural knowledge in constructing global features. 2) The lower layers of the pre-trained backbone from BEV map generation are shared for visual and structural streams in VPR, facilitating the learning of fine-grained local features in the visual stream. 3) The complementary viusal features and structural features can jointly enhance VPR performance. Our BEVPR framework enables consistent performance improvements over several popular camera-based VPR aggregation modules when integrating them. The experiments on our collected VPR-NuScenes dataset demonstrate an absolute gain of 2.47% on Recall@1 for the strong Conv-AP baseline to achieve the best performance in our setting, and notably, a 18.06% gain on the hard set. The code and VPR-NuScenes dataset will be available at https://github.com/FudongGe/BEV2PR.
@@ I INTRODUCTION
When locating, the human brain typically relies on the sights observed by eyes, including the colors of scenes, the textures of objects’ surfaces, and the spatial structure perception learned through years of experience. Therefore, it’s possible to construct a robust representation with both appearance and structural properties solely based on the camera. After reviewing the development progress of VPR technology, we find that the research based on both appearance and structure is a hotter field compared to studies solely based on appearance or structure [5, 6, 7]. As depicted in Fig. 1, existing methods based on both appearance and structure can be divided into approaches using both camera and LiDAR sensors [5, 6, 7], i.e., (b), and camera-based approaches [8, 9, 10], i.e., (a), (c), (d) and (e). Furthermore, the structural information sources of camera-based methods include reconstructed 3D point clouds, 2D segmentation images, 2D depth images, thus (c) . (d) . (e) is also a focus worth analyzing.
@@
Upon examining the current methods, it is noted that each possesses its own advantages and disadvantages. The methods based on both camera and LiDAR in Fig. 1(b), which utilize the complementarity of multimodal raw information, face an issue involving higher procurement costs and maintenance expenses, as well as the calibration and alignment of different modal data [5, 6, 7]. Vehicles not equipped with LiDAR cannot deploy such algorithms. The SfM (Structure-from-Motion)-based methods with pseudo-3D point clouds reconstructed from RGB images, i.e., Fig. 1(d), reduces the costs of traditional approaches. However, this reconstruction process is relatively complex, and the data required for inference are more challenging to handle than conventional pixel images [9]. Moreover, this method highly depends on the quality and coverage of the images. If the images are of poor quality (for example, blurred due to rain), or if there is insufficient coverage between different images, it may lead to low-quality reconstructed point clouds, thereby impacting the accuracy. Beneficial from current segmentation models [11], SEG-based methods in Fig. 1(c) use segmentation images to enhance structural knowledge of global features. However, segmentation images are still in the x-y 2D plane without the implicit structure - depth, resulting in the loss of key information. Fig. 1(e) provides depth information, while depth images contain harmful noise from dynamic objects, and lack explicit spatial relationships between different objects. The limitations of the current methods raise a question: Could we integrate explicit depth and spatial relationships as well as RGB information into global features using images as input during inference?
@@
To tackle this issue, we propose our solution - RGB and BEV fusion, inspired by the widespread application of BEV representations in 3D perception tasks, where BEV exhibits outstanding performance in clearly depicting the relative positions and shapes of objects. The challenge lies in how to integrate BEV features into global features and how to leverage the knowledge in structural stream to enhance visual stream. Therefore, we propose a new architecture, BEVPR, which is designed to simultaneously construct semantic map to model spatial relationships in the BEV frame, and generate a composite descriptor with both visual cues and spatial awareness for VPR. Specifically, we first pre-train a BEV generation model to extract BEV features as a more explicit source of structural knowledge. And we copy the modules of the BEV model as part of the structural stream and freeze them to introduce BEV features into VPR, then crop its backbone into two parts, with the former serving as the bottom backbone shared with the visual stream and the latter copied as the sub-backbone of visual stream and then unfrozen. After that, any popular aggregation module for RGB global features is inserted into the visual stream, followed by a fusion operation. In particular, our framework only uses images as input, without involving other sensors or complex training processes, making it generalized for a wide range of autonomous vehicles.
@@
Our main contributions can be highlighted as follows: 1) Data Module: We introduce VPR-NuScenes to distinguish between simple and challenging scenes to understand the respective strengths of visual and structural representations. 2) Architecture: We propose BEVPR to simultaneously construct semantic map to model spatial relationships in the BEV frame, and generate a composite descriptor with both visual cues and spatial awareness for VPR. 3) Experimental Results: We conduct extensive experiments to evaluate the improvement effect of our framework on different appearance-based methods, particularly in scenes characterized by significant appearance variations.
@@ II RELATED WORK
The related work in this paper broadly involves appearance-based place recognition as well as place recognition based on both appearance and structure. The former focus on studying the semantic and texture information of RGB images, while the latter introduces structural information on the basis of the former.
II-A Place Recognition based on Appearance
The appearance-based place recognition concentrates on constructing better image representations including global or local descriptors for retrieval, where global descriptors can be generated through direct extraction or by aggregating local descriptors [10]. As a representative of aggregation technology, NetVLAD [12] is a trainable variant of VLAD, softly allocating local features to a set of learned clusters. Based on it, many variants have been inspired, such as SPE-NetVLAD [13] and Gated NetVLAD [14].
@@ II-A Place Recognition based on Appearance
II-B Place Recognition based on both Appearance and Structure
Place Recognition Relying on LiDAR. Considering the robustness of structural features in some environments, scholars have attempted to introduce the technology of fusing lidar and cameras into place recognition. PIC-Net [19] by Lu et al. employs global channel attention to enhance the interaction between image and point cloud features, along with a spatial attention-based VLAD to select the discriminative points and pixels. Zhou et al. develop LCPR [7], a multi-scale network leveraging self-attention mechanism to correlate panoramic features across different modalities. OneShot [20] by Ratz et al. projects the segments from point clouds onto images to facilitate feature extraction through 2D and 3D CNN. CORAL [21] by Pan et al. first builds an elevation image from lidar scans, and then augments it with projected visual features to generate a fusion descriptor in the BEV frame.
Place Recognition Independent of LiDAR. Another path relies solely on cameras. Oertel et al. [9] use structural features extracted from image sequences through vision-based SfM to augment VPR. Shen et al. [10] try to leverage segmentation images obtained from a pre-trained segmentation model to augment the structural comprehension embedded within global representations, achieving state-of-the-art performance. Based on knowledge transfer and adversarial learning, Qin et al. [22] propose a structure-aware feature disentanglement network, called SFDNet, where probabilistic knowledge transfer is employed to transfer knowledge obtained from the Canny edge detector to the structure encoder. DASGIL [8] by Hu et al. uses multi-task architecture with a single shared encoder to create global representation.
Nevertheless, these previous works fail to simultaneously meet the requirements of low cost as well as explicit spatial relationships within global features. In this paper, we for the first time try to enhance VPR through a late-fusion approach of RGB features and BEV structural features.
@@ III METHODOLOGY
Considering that BEV representation contains enriched and explicit structural knowledge, we attempt to leverage BEV features to infuse structural information into global retrieval. The core concept behind BEVPR focuses on the static features in BEV images and the explicit spatial relationships between different objects, as well as the single-input but dual-modal processing method. BEVPR involves two training stages: Camera-to-BEV training and VPR training, which is shown how to work in Fig. 2.
@@ III-A Overview
Given an input RGB image, a BEV generation model is pre-trained in the first training stage, by which BEV features can be obtained to assist in VPR. In the second training stage, the BEV model is frozen and its backbone is cropped into two sequential parts. Finally, we insert any global feature aggregation into visual stream, followed by a fusion operation to obtain a composite descriptor.
@@ III-B BEV Map Generation
Semantic Class Selection. In various BEV images, there are different numbers of semantic instances, making the selection of appropriate types and quantities significant. A BEV map is represented by a tensor, where represents the number of semantic classes, and corresponds to the area in front of the ego vehicle. Within this tensor, regions belonging to the class are marked with positive values in the corresponding channel, while all other channels are set to zero. And, in the context of VPR, we improve model convergence and accuracy by only using static classes as the target to train the BEV generation model. This technique draws inspiration from the fact that each semantic class plays a different role in VPR, and further human brain’s attention mechanism, which concentrates on iconic static elements or important regions, such as roads and buildings, while ignoring dynamic ones similar to noise.
Camera-to-BEV Model. In this section, we establish a BEV generation model with only a front-view image as input. First, an image is embedded by a 2D CNN to get the feature map. And then, we transform the feature map from 2D to 3D using a similar method with [23], that is 1) predict grid-wise depth distribution with equal spacing on 2D features, 2) ‘lift’ the 2D features to voxel space based on depth, obtaining a pseudo point cloud features. Finally, a pooling operation similar to LiDAR-based methods is implemented to flatten 3D point cloud features into 2D BEV features.
@@ III-C Structural Knowledge Extraction
Shared Bottom Backbone. Due to the insensitivity of global features to small targets in the raw visual stream, we adopt a shared bottom backbone instead of the architecture of separately processing input images in visual and structural streams. To specify, we crop the backbone of BEV generation model into two sequential parts, with the former serving as a bottom backbone shared with the visual stream, so as to prompt visual stream to learn fine-grained local features generated by the BEV semantic segmentation, after which, a specific sub-backbone for the visual stream is used to capture important features for VPR. The experimental results in Section IV confirm its effectiveness.
@@
III-D Feature Aggregation
As shown in Fig. 2, there are two feature aggregation modules in visual and structural streams, one of which in the structural stream is set to use GeM. In our work, we evaluate multiple methods to verify the improvement effects of BEV features on them. The feature map in visual or structural stream is denoted as .
@@ III-D Feature Aggregation
where .
• NetVLAD [12]: As a representative of aggregation technology, it’s a trainable variant of VLAD, softly allocating local features to a set of learned clusters, defined as
@@
where , and are sets of trainable parameters for each cluster, represents the total number of features.
• GeM [15]: This method implements a parametric generalized-mean mechanism, which can assign a shared parameter per feature map. Owing to the differentiable nature of the pooling operation, the parameters can be learned as part of the back-propagation. GeM is defined as
where , represents the learnable parameter.
• Conv-AP [18]: This is a fully convolutional feature aggregation technique, which performs channel-wise pooling on the feature maps followed by spatial-wise adaptive pooling, making the architecture convolutional and the output dimensionality configurable. Conv-AP is defined as
where represents the number of spatial sub-regions.
• EigenPlaces [26]: As a variant that builds on GeM aggregator, EigenPlaces utilizes a GeM pooling and a fully connected layer with output dimension 512. It’s defined as
• MixVPR [27]: This work focuses on an isotropic all-MLP architecture, and utilizes feature maps extracted from a pre-trained backbone as an aggregation of global descriptors to integrate the global interaction among elements within each feature map through a cascade of feature mixing. First, is considered as a set of 2D features, i.e., , , followed by a flattening operation, resulting in flattened feature maps , where . Then, it’s fed into Feature Mixer blocks, i.e.,
@@ III-E Feature Fusion
Here, we use a weighted concatenation with two learnable scalar weights and rescaling each of the vectors, and the final descriptor is given by
where and represent the feature vectors of visual stream and structural stream respectively.
@@ III-F Network Training
We train our compound network in a two-stage manner instead of the extremely slow single-stage method. The loss function in the first stage fully refers to existing work [23, 28]. Then, following [12], we adopt the triplet margin loss to train the second-stage network. To specify, each training iteration involves a mini-batch consisting of one anchor, multiple anchor’s positive matches and negative matches. More details on the partitioning of positive and negative samples can be found in Section IV.A. We define as the function mapping input to its descriptor, and the loss function of each descriptor is defined as
where represent the type of three descriptors, and represent the number of positive and negative samples in each mini-batch, represent a tuple loss, i.e.,
where computes the Euclidean distance of two descriptors, margin is a constant.
Moreover, it has been reported in [5] that the network tends to overfit in image domain during training. Therefore, we use a multi-head loss function, denoted as

### v2

# BEV2PR: BEV-Enhanced Visual Place Recognition with Structural Cues

@@ Abstract
In this paper, we propose a new image-based visual place recognition (VPR) framework by exploiting the structural cues in bird’s-eye view (BEV) from a single monocular camera. The motivation arises from two key observations about place recognition methods based on both appearance and structure: 1) For the methods relying on LiDAR sensors, the integration of LiDAR in robotic systems has led to increased expenses, while the alignment of data between different sensors is also a major challenge. 2) Other image-/camera-based methods, involving integrating RGB images and their derived variants (e.g., pseudo depth images, pseudo 3D point clouds), exhibit several limitations, such as the failure to effectively exploit the explicit spatial relationships between different objects. To tackle the above issues, we design a new BEV-enhanced VPR framework, namely BEV2PR, generating a composite descriptor with both visual cues and spatial awareness based on a single camera. The key points lie in: 1) We use BEV features as an explicit source of structural knowledge in constructing global features. 2) The lower layers of the pre-trained backbone from BEV generation are shared for visual and structural streams in VPR, facilitating the learning of fine-grained local features in the visual stream. 3) The complementary visual and structural features can jointly enhance VPR performance. Our BEV2PR framework enables consistent performance improvements over several popular aggregation modules for RGB global features. The experiments on our collected VPR-NuScenes dataset demonstrate an absolute gain of 2.47% on Recall@1 for the strong Conv-AP baseline to achieve the best performance in our setting, and notably, a 18.06% gain on the hard set. The code and dataset will be available at https://github.com/FudongGe/BEV2PR.
@@ I INTRODUCTION
After reviewing the progress in place recognition methods, we find that the research based on both appearance and structure is a hotter field [5, 6, 7]. Despite the primary focus of this paper being on VPR tasks, we also introduce some methods related to LiDAR to emphasize the importance of structural features. As depicted in Fig. 1, existing methods based on both appearance and structure can be divided into approaches using both camera and LiDAR sensors [5, 6, 7], i.e., (b), and camera-based approaches [8, 9, 10], i.e., (a), (c), (d) and (e). Furthermore, the structural information sources of camera-based methods include 2D segmentation images, reconstructed 3D point clouds, 2D depth images, thus (c) . (d) . (e) is also a focus worth analyzing.
@@
Upon examining the current methods, it is noted that each possesses its own advantages and disadvantages. The methods based on both camera and LiDAR in Fig. 1(b), which utilize the complementarity of multimodal raw information, face issues involving expensive equipments, as well as the calibration and alignment of different modal data [5, 6, 7]. Vehicles not equipped with LiDAR cannot deploy such algorithms. The SfM (Structure-from-Motion)-based methods with pseudo 3D point clouds reconstructed from RGB images, i.e., Fig. 1(d), avoids the expense of using LiDAR. However, this reconstruction process is relatively complex, and the data required for inference are more challenging to handle than conventional pixel images [9]. Beneficial from current segmentation models [11], SEG-based methods in Fig. 1(c) use segmentation images to enhance structural knowledge of global features. However, segmentation images are still in the x-y 2D plane without the implicit structure - depth, resulting in the loss of key information. Fig. 1(e) provides depth information, while depth images contain harmful noise from dynamic objects, and lack explicit spatial relationships between different objects. The limitations of the current methods raise a question: Could we integrate explicit depth and spatial relationships as well as RGB information into global features using images as input during inference?
@@
To tackle this issue, we propose our solution - RGB and BEV fusion, inspired by the widespread application of bird’s-eye view (BEV) representations in 3D perception tasks, where BEV exhibits outstanding performance in clearly depicting the relative positions. The challenge lies in how to integrate BEV features into global features and leverage knowledge from the structural stream to enhance the visual stream. Therefore, we propose BEV2PR, a new architecture simultaneously constructing semantic map to model spatial relationships in the BEV frame, and generating a composite descriptor with both visual cues and spatial awareness for VPR. Specifically, we first pre-train a BEV generation model to extract BEV features as a more explicit source of structural knowledge. And we copy the modules of this model as part of the structural stream and freeze them to introduce BEV features into VPR, then crop its backbone into two parts, with the former serving as the bottom backbone shared with visual stream and the latter copied as the sub-backbone of visual stream and then unfrozen. The details are shown in Fig. 2. After that, any aggregation module for RGB global features is inserted into the visual stream, followed by a fusion operation. In particular, our framework only uses an image as input, without involving other sensors or complex training processes, making it generalized for a wide range of autonomous vehicles.
@@
Our main contributions can be highlighted as follows: 1) Data Module: We introduce VPR-NuScenes to distinguish between simple and challenging scenes to understand the respective strengths of visual and structural representations. 2) Architecture: We propose BEV2PR to simultaneously construct semantic map to model spatial relationships in the BEV frame, and generate a composite descriptor with both visual cues and spatial awareness for VPR. 3) Experimental Results: We conduct extensive experiments to evaluate the improvement effect of our framework on different appearance-based methods, particularly in scenes characterized by significant appearance variations.
@@ II RELATED WORK
The related work involves appearance-based place recognition as well as place recognition based on both appearance and structure. The former focuses on studying the semantic and texture information of RGB images, while the latter introduces structural information on the basis of the former.
II-A Place Recognition with Appearance
The appearance-based place recognition concentrates on constructing better image representations including global or local descriptors for retrieval, where global descriptors can be generated through direct extraction or by aggregating local descriptors [10]. As a representative of aggregation algorithm, NetVLAD [12] is a trainable variant of VLAD, softly allocating local features to a set of learned clusters. Based on it, many variants have been inspired, such as SPE-NetVLAD [13] and Gated NetVLAD [14].
@@ II-A Place Recognition with Appearance
II-B Place Recognition with both Appearance and Structure
Place Recognition Relying on LiDAR. Considering the robustness of structural features in some environments, scholars have attempted to introduce fusion technologies of point cloud and image into place recognition. PIC-Net [19] by Lu et al. employs global channel attention to enhance the interaction between point cloud and image features, along with a spatial attention-based VLAD to select the discriminative points and pixels. Zhou et al. develop LCPR [7], a multi-scale network leveraging self-attention mechanism to correlate panoramic features across different modalities. OneShot [20] by Ratz et al. projects the segments from point clouds onto images to facilitate feature extraction through 2D and 3D CNN. CORAL [21] by Pan et al. first builds an elevation image from LiDAR scans, and then augments it with projected visual features to generate a fusion descriptor.
Place Recognition Independent of LiDAR. Another path relies solely on cameras. Oertel et al. [9] use structural features extracted from image sequences through vision-based SfM to augment VPR. Shen et al. [10] leverage segmentation images obtained from a pre-trained segmentation model to augment the structural comprehension embedded within global representations, achieving state-of-the-art performance. Based on knowledge transfer and adversarial learning, Qin et al. [22] propose a structure-aware feature disentanglement network, called SFDNet, which leverages probabilistic knowledge transfer to transmit features from Canny edge detector to structure encoder. DASGIL [8] by Hu et al. uses a multi-task architecture to integrate geometric and semantic knowledge into multi-scale global representations.
In this paper, we construct semantic map to model spatial relationships within the BEV frame, and for the first time, adopt a late-fusion approach of RGB and BEV features to enhance VPR, benefitting from the explicit spatial relationship modeling between different objects in BEV generation.
@@ III METHODOLOGY
Considering the rich and explicit structural knowledge inherent in BEV representations, we leverage BEV features to infuse structural information into global retrieval. The core concept behind BEV2PR focuses on the static features in BEV images and the explicit spatial relationships between different objects, as well as the single-input but dual-modal processing method. BEV2PR involves two training stages: Camera-to-BEV training and VPR training shown in Fig. 2.
@@ III-A Overview
Given an input RGB image, a BEV generation model is pre-trained in the first training stage, by which BEV features can be obtained to assist in VPR. In the second training stage, the BEV model is frozen and its backbone is cropped into two sequential parts. Finally, we insert any global feature aggregation module into the visual stream, followed by a fusion operation to obtain a composite descriptor.
@@ III-B BEV Map Generation
Semantic Class Selection. In various BEV images, there are different numbers of semantic instances, making the selection of appropriate types and quantities significant. A BEV map is represented by a tensor, where represents the number of semantic classes, and corresponds to the area in front of the ego vehicle. Within this tensor, regions belonging to the class are marked with positive values in the corresponding channel, while those in other channels are set to zero. And, in the context of VPR, we improve model convergence and accuracy by only using static classes as the target to train the BEV generation model. This technique draws inspiration from the fact that each semantic class plays a different role in VPR, and further human brain’s attention mechanism, which concentrates on iconic static elements or important regions, such as roads and buildings, while ignoring dynamic ones similar to noise.
Camera-to-BEV Model. In this section, we establish a BEV generation model with only a front-view image as input. First, an image is embedded by a 2D CNN to get its feature map. And then, we transform the feature map from 2D to 3D using a similar method with [23], that is 1) predict grid-wise depth distribution with equal spacing on 2D features, 2) ‘lift’ the 2D features to voxel space based on depth, obtaining pseudo point cloud features. Finally, a pooling operation similar to LiDAR-based methods is implemented to flatten 3D point cloud features into 2D BEV features.
@@ III-C Structural Knowledge Extraction
Shared Bottom Backbone. Due to the insensitivity of global features to small targets in the raw visual stream, we adopt a shared bottom backbone instead of the architecture of separately processing input images in visual and structural streams. To specify, we crop the backbone of BEV generation model into two sequential parts, with the former serving as a bottom backbone shared with the visual stream, so as to prompt visual stream to learn fine-grained local features generated by the BEV semantic segmentation, meanwhile, the latter is copied as the specific sub-backbone of visual stream and unfrozen to capture important features for VPR.
@@
III-D Preliminary Feature Aggregation
In our work, we evaluate multiple methods to verify the improvement effect of BEV features on them. The feature in the visual stream is denoted as , where represents the number of features maps. And is defined as the 2D feature map of dimensions, where . Let be the output representation.
@@ III-D Preliminary Feature Aggregation
• NetVLAD [12]: As a representative of aggregation algorithm, it is a trainable variant of VLAD, softly allocating local features to a set of learned clusters, defined as
@@
where represents the descriptor, represents the cluster center, and are sets of trainable parameters for each cluster, represents the total number of features. Refer to [12] for more details.
• GeM [15]: This method implements a parametric generalized-mean mechanism, which can assign a shared parameter per feature map. Owing to the differentiable nature of pooling operation, the parameters can be learned as part of the back-propagation. The global feature is defined the same as Eq. 1, but the local feature is denoted as
where represents the learnable parameter.
• Conv-AP [18]: This is a fully convolutional feature aggregation technique, which implements channel-wise pooling on feature maps, coupled with spatial-wise adaptive pooling, allowing for significant configurability in output dimensionality. Conv-AP is defined as
where represents the adaptive average pooling, represents the number of spatial sub-regions, represents the convolution.
• EigenPlaces [26]: As a variant that builds on GeM aggregator, EigenPlaces utilizes a GeM pooling and a fully connected (FC) layer with output dimension 512, defined as
• MixVPR [27]: This work focuses on an isotropic all-MLP architecture, and utilizes feature maps extracted from a pre-trained backbone as an aggregation of global descriptors to integrate the global interaction among elements within each feature map through a cascade of feature mixing. First, is considered as a set of 2D features, i.e., , followed by a flattening operation, resulting in feature maps , where . Then, it is fed into Feature Mixer blocks, i.e.,
@@
Finally, is flattened and -normalized as usually done in VPR [12], still recorded as for convenience.
@@ III-E Feature Fusion
Here, we use a weighted concatenation with two learnable scalar weights and to rescale each descriptor in two streams, and the feature aggregation module in the structural stream is set to use GeM with the output representation denoted as . Then the final descriptor is given by
@@ III-F Network Training
We train our compound network in a two-stage manner. The loss function in the first stage fully refers to existing work [23, 28]. For the second stage, each training iteration involves a mini-batch consisting of one query sample as well as multiple positive matches and negative matches for it. Following [12], we adopt the triplet margin loss for each tuple, i.e.,
where and represent the positive sample and the negative sample for in a mini-batch, margin is a constant, computes the Euclidean distance of two descriptors, is defined as the function mapping input to its descriptor. Then, the loss function of each descriptor is defined as
where and represent the number of positive and negative samples in each mini-batch, represent the type of three descriptors, i.e., fusion descriptor loss , visual descriptor loss , structural descriptor loss . Considering that the network tends to overfit in one single domain during training [5], we use a multi-head loss function for final back-propagation, denoted as
@@
IV EXPERIMENTS
IV-A Dataset
Partitioning of Positive and Negative Samples. Given the spatial proximity, the limited variations in camera angles and the minimal changes in appearance between consecutive frames, treating consecutive frames as positive samples proves excessively facile and lacks practical significance. Consequently, images within the same scene are not regarded as positive samples to one another. The selection of negative samples is straightforward: for each query sample, samples other than its positive samples and consecutive frames are defined as negative samples.